If it takes 12 2/3 gallons of gas to fill a car's tank, how much does it take to fill 3/5 of the tank with gas?
Answer by Cromlix(4381)

Hi there,
12 2/3 gallons = a full tank
38/3 gallons = a full tank
3/5 x 38/3 = 3/5 full tank.
= 114/15 = 7 3/5 gallons
So, it takes 7 3/5 gallons to
fill 3/5 of the tank.
Hope this helps:-)